OpenStudy (anguyennn):

how to simplify 2√4-√9?

jimthompson5910 (jim_thompson5910):

What is the square root of 4 equal to?

OpenStudy (anguyennn):

2

jimthompson5910 (jim_thompson5910):

correct What is the square root of 9 equal to?

OpenStudy (anguyennn):

3

OpenStudy (anguyennn):

and you multiply 2x2 and minus 3? correct?

jimthompson5910 (jim_thompson5910):

so you really have `2*2-3` evaluate that to get ??

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anguyennn):

oh ok!! thank you but what about when its √2?

jimthompson5910 (jim_thompson5910):

taking the square root of 2 doesn't give a whole number like the square root of 4 does

OpenStudy (anguyennn):

like 3√2-5√2-2√2

jimthompson5910 (jim_thompson5910):

I don't see a square root of 2 in your problem though

jimthompson5910 (jim_thompson5910):

Oh I see now

jimthompson5910 (jim_thompson5910):

replace every copy of "square root of 2" with x so you'll have 3x-5x-2x simplify that to get ???

OpenStudy (anguyennn):

-4x

OpenStudy (anguyennn):

but how come you replaced it with an x?

jimthompson5910 (jim_thompson5910):

well you don't need to, but you can think of it like that basically you're combining like terms. In this case, all 3 terms are like terms because they have a root 2 in them

OpenStudy (anguyennn):

yes

jimthompson5910 (jim_thompson5910):

3x-5x-2x = -4x so you now replace the x with square root of 2 so the final answer is \(\Large -4\sqrt{2}\)

OpenStudy (anguyennn):

i am still confused because for other questions it would have different numbers and i don't know how to do them

jimthompson5910 (jim_thompson5910):

like what for example?

OpenStudy (anguyennn):

√18+√50-√8

jimthompson5910 (jim_thompson5910):

have you tried to simplify each root?

OpenStudy (anguyennn):

dividing them all by 2?

jimthompson5910 (jim_thompson5910):

think of all of the factors of 18 which factors are perfect squares?

OpenStudy (anguyennn):

I am not sure

OpenStudy (anguyennn):

sorry

jimthompson5910 (jim_thompson5910):

factors of 18 1,2,3,6,9,18

OpenStudy (anguyennn):

ok

jimthompson5910 (jim_thompson5910):

which factor is a perfect square?

OpenStudy (anguyennn):

9

jimthompson5910 (jim_thompson5910):

so we can say this \[\Large \sqrt{18} = \sqrt{9*2}\] \[\Large \sqrt{18} = \sqrt{9}*\sqrt{2}\] \[\Large \sqrt{18} = 3*\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

I factored 18 into 9*2 then I used the rule \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]

OpenStudy (anguyennn):

yes

jimthompson5910 (jim_thompson5910):

now let's do 50 what are the factors of 50?

OpenStudy (anguyennn):

1, 2, 5, 10, 25

OpenStudy (anguyennn):

and it would be 25 that is the perfect square

jimthompson5910 (jim_thompson5910):

yes so 50 = 25*2

jimthompson5910 (jim_thompson5910):

what would \(\Large \sqrt{50}\) simplify to?

OpenStudy (anguyennn):

√25x2

jimthompson5910 (jim_thompson5910):

then break it up using the rule I posted

OpenStudy (anguyennn):

√50= √25x2 =√25x √2 5√2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now onto 8 the factors of 8 are ???

OpenStudy (anguyennn):

1,2,4,8

jimthompson5910 (jim_thompson5910):

which is a perfect square?

OpenStudy (anguyennn):

√4

jimthompson5910 (jim_thompson5910):

so we can say 8 = 4*2

jimthompson5910 (jim_thompson5910):

then use that rule

OpenStudy (anguyennn):

√8= √4x2 =√4x√2 2√2

jimthompson5910 (jim_thompson5910):

so, \[\Large \sqrt{18}+\sqrt{50}-\sqrt{8}\] is the same as \[\Large 3\sqrt{2}+5\sqrt{2}-2\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

we can replace all the square root of 2 terms with x \[\Large 3\color{red}{\sqrt{2}}+5\color{red}{\sqrt{2}}-2\color{red}{\sqrt{2}}\] \[\Large 3\color{red}{x}+5\color{red}{x}-2\color{red}{x}\]

OpenStudy (anguyennn):

correct

OpenStudy (anguyennn):

is that the answer?

jimthompson5910 (jim_thompson5910):

not yet

OpenStudy (anguyennn):

would you do 3+5-2 now?

OpenStudy (anguyennn):

and get 6√2?

jimthompson5910 (jim_thompson5910):

very good, that's your final answer

OpenStudy (anguyennn):

oh ok thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome